OpenStudy (anonymous):

x-3 (underneath the radical) = 2 a)x= -1 b) x = 1 c) x = 5 d) x = 7 ?

OpenStudy (anonymous):

\[\sqrt{x-3}=2 *\]

OpenStudy (anonymous):

square both sides and u get x-3=4. so x=7.

OpenStudy (anonymous):

but when yu square 3 and then square 2 9 - 4= 5 , so wouldnt it be c?

OpenStudy (anonymous):

2nd question . . ^